# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ 3.3 External validity: unknown deployment shift
However, the same properties of multi-accuracy also imply robustness to external shift. In this subsection, we indeed suppose Assumption 1, unconfoundedness. Recall that our goal was to control the predictive bias on a reference covariate distribution $Q_{X}$, potentially unknown, $\lvert\operatorname{E}_{Q}[\left(\hat{\tau}(X)-\tau(X)\right)]\rvert$. Note that each of $\mu_{1},\mu_{0}$ are learned on a treatment conditional distribution, so we have that the valid likelihood ratio, which we denote $w_{t}(x)$, is defined as:
@@
Suppose Assumptions 1 and 2. Let $\mathcal{C}$ denote a test function class for subgroup membership and $\mathcal{H}$ a test function class for likelihood ratios. Then multi-accuracy of the T-learner CATE estimate by running Algorithm 1 implies that, for all reference covariate distributions $Q$ such that the likelihood ratios $w_{1},w_{0}\in\mathcal{H},$
$\forall c\in\mathcal{C},\;\lvert\operatorname{E}_{Q}[\{\tilde{\tau}(X)-(Y(1)-Y(0))\}c(X)]\rvert\leq 2\alpha$
@@
Observe that although similar arguments apply, obtaining conditional guarantees for CATE estimation requires a richer test function class than for universal adaptability of the ATE alone. This illustrates that the case of learning CATE is indeed statistically harder than that of “universal adaptability” of the ATE that was studied in Kim et al. [2022]. For CATE estimation, we need to choose a richer auditor function class than we would for ATE estimation.
3.3.1 Extension to CATE pseudo-outcome regression
A natural question given our work on the T-learner is whether we can provide similar guarantees for an estimation-improved CATE learner, since the T-learner generally does not enjoy any improved estimation properties in causal inference, whereas causal inference and machine learning has developed many improved orthogonal/semiparametrically efficient procedures such as (but not limiited to) the R-learner [Nie and Wager, 2020], DR-learner [Kennedy, 2023], or other machine-learning adaptations [Wager and Athey, 2018, Shalit et al., 2017].
Namely, some CATE estimation procedures give a pseudo-outcome $\psi(O;e,\mu)$, where $O$ denotes data tuples, i.e. $O=(X,T,Y)$, such that $\operatorname{E}[\psi(O;e,\mu)\mid X]=\tau(X)$. (It is designated as a pseudo-outcome because regressing upon it identifies the CATE or functional of interest, although it is not exactly outcome itself). One such pseudo-outcome is the doubly-robust score. Pseudo-outcome regression of it as a CATE estimator was recently studied in Semenova and Chernozhukov [2021], Kennedy [2020].
@@ 3.3.1 Extension to CATE pseudo-outcome regression
Regressing upon pseudo-outcomes with favorable properties such as orthogonal moment conditions therefore confers such favorable properties to the estimated functional, such as improved statistical rates of convergence. Our arguments for external validity can naturally be extended for pseudo-outcome based CATE regression, so long as the pseudo-outcome’s conditional expectation is the CATE function.
We multi-calibrate the pseudo-outcome regression step. That is, we learn $\tilde{\tau}$ such that:
@@
Multi-accurate DR-learner
We give the algorithm for obtaining a multi-accurate DR-learner estimate in Algorithm 2. To summarize: we do need four folds of data $\left(\mathcal{D}_{1a},\mathcal{D}_{1b},\mathcal{D}_{2},\mathcal{D}_{3}\right)$; the first three for sample-splitting of the nuisance estimates and pseudo-outcome evaluation and the last for validation/calibration for MCBoost. Estimate the nuisance functions on the first two folds $\mathcal{D}_{1a},\mathcal{D}_{1b}$ and on $\mathcal{D}_{2},$ evaluate the pseudo-outcome value $\hat{\varphi}(O;\hat{e},\hat{\mu})$ and regress $\hat{\tau}(x)=\hat{\mathbb{E}}_{n}\{\hat{\varphi}(O;\hat{e},\hat{\mu})\mid X=x\}.$ Finally, we conduct post-processing via multi-accurate learning upon the DR-learner estimate $\hat{\tau}$, to obtain a multi-accurate $\tilde{\tau}$.
@@
$\operatorname{E}[\{\hat{\varphi}(\hat{e},\hat{\mu})-\tilde{\tau}(X)\}f(x)]\leq\epsilon,\forall f\in\mathcal{F}$
5:Cross-fitting (optional)222 omitted for brevity because this may have to be further modified: original text reads, Repeat Step 1-2 twice, first using $\left(D_{1b}^{n},D_{2}^{n}\right)$ for nuisance training and $D_{1a}^{n}$ as the test sample, and then using $\left(D_{1a}^{n},D_{2}^{n}\right)$ for training and $D_{1b}^{n}$ as the test sample. Use the average of the resulting three estimators as a final estimate of $\tau$.
@@
Suppose that $\tilde{\mu}_{t}(x)\leftarrow\textrm{MCBoost}(\hat{\mu}_{t},\alpha,\mathcal{F},\mathcal{D}_{post})$, over auditor function class $\mathcal{F}$ such that ¯$\overline{\mathcal{F}}_{t}\subseteq\mathcal{F}$, where ¯$\overline{\mathcal{F}}_{t}=\left\{1,\frac{\mathbb{I}[T=t]}{e_{t}(X)}\right\}\times\mathcal{C}\times\mathcal{H}$. That is, $\tilde{\mu}_{1}-\tilde{\mu}_{0}$ comprise an $\alpha$-multi-accurate $T$ learner. Then
$\Big{|}{\max_{ch\in\mathcal{C}\times\mathcal{H}}\operatorname{E}[\{{\varphi}(e,\tilde{\mu})-\tau(X)\}c(x)h(x)]-\max_{ch\in\mathcal{C}\times\mathcal{H}}\operatorname{E}[\{\tilde{\tau}-\tau(X)\}c(x)h(x)]}\Big{|}\leq 2\alpha$
That is, the multi-accurate T-learner $\tilde{\mu}_{1}-\tilde{\mu}_{0}$ is, up to $2\alpha$ additive approximation error, a multi-accurate DR-learner with outcome model $\tilde{\mu}$, post-processed over the function class $\mathcal{C}\times\mathcal{H}.$ [az: fixup statement]
@@
$\max_{c\times h\in\{\mathcal{C}\times\mathcal{H}\}}\lvert E[\{\{(Y-\tilde{\mu}_{1}(X))\frac{\mathbb{I}[T=1]}{e_{1}(X)}+(Y-\tilde{\mu}_{0}(X))\frac{\mathbb{I}[T=0]}{e_{0}(X)}\}+\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)\}-\tau(X)\}c(X)h(X)]\rvert$
$\qquad-\max_{c\times h\in\{\mathcal{C}\times\mathcal{H}\}}\lvert E[\{\{\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)\}-\tau(X)\}c(X)h(X)]\rvert$
@@
Interestingly, concurrent with the preparation of this work, [Bruns-Smith et al., 2023] study augmented balancing weights and find a certain target-independent property of the augmented estimator related to the universal adaptability of [Kim et al., 2022]. Studying connections further would be an interesting direction for future work.
3.4 Observational and Randomized data (Setting 2)
(Meta)-Algorithm
Identification of CATE
Identification for the CATE follows by interpreting the auditing functions $c(X)\in\mathcal{C}$ as subpopulations. Achieving multi-accuracy on the RCT data hence identifies the CATE. That is, multi-accuracy assures us that
$\hat{\mu}_{t}(x)\leftarrow\arg\min_{g\in\mathcal{G}}\operatorname{E}_{\text{obs}}[(g-Y)^{2}\mid T=t],\text{ for }t\in\{0,1\}$
4:Return $\tilde{\tau}(x)=\tilde{\mu}_{1}(x)-\tilde{\mu}_{0}(x)$
Identification of target-independent CATE
In complete analogy to the external shift setting, changing our interpretation of the target functions allows us to infer robustness to external shifts. Multi-accuracy ensures that, for all reference covariate distributions $Q_{X}$ such that the likelihood ratios $w_{t}(x)\in\mathcal{H},\;t\in\{0,1\}$, running Algorithm 3 with auditor function class $\mathcal{F}=\mathcal{C}\times\mathcal{H}$ results in a multi-accurate and deployment-shift robust CATE estimate:
$\forall c\in\mathcal{C},\;\lvert\operatorname{E}_{Q}[\{\tilde{\tau}(X)-(Y(1)-Y(0))\}c(X)]\rvert\leq 2\alpha.$
@@ 4.2 WHI data application
We next present a case study that draws on data from the Women’s Health Initiative (WHI) studies [Machens and Schmidt‐Gollwitzer, 2003]. The WHI includes a large observational study and clinical trial data to investigate the effectiveness of hormone replacement therapy (HRT) in preventing the onset of chronic diseases. As the observational study has been suspected to suffer from various (unobserved) confounding phenomena [Kallus and Zhou, 2018], we study how utilizing both data sources in combination via multi-accuracy boosting compares to learning CATE from the observational or clinical trial data only. We focus on the effect of HRT treatment on systolic blood pressure and use age and ethnicity as covariates. Implementation details and results with an extended set of covariates is presented in the supplementary material D.3.
@@ 5 Related work: further discussion
A popular approach for handling unknown shifts is to enforce robustness against a family of covariate shifts (e.g. unknown shifts parametrized by unknown covariate shift functions) [Liu and Ziebart, 2014, Wen et al., 2014, Chen et al., 2016]. The goal is to find a robust hypothesis that maximizes the worst-case prediction risk (for example, squared error) evaluated with respect to unknown shifts within some class of covariate shifts. Parametrizations include distributionally robust optimization or linear basis functions. The work of [Greenfeld and Shalit, 2020] is motivated differently and penalizes with a Hilbert-Schmidt Independence Criterion loss; they show this implies some robustness to covariate shift. While most of this work is in the generic prediction setting, recent work also assesses ATE under covariate shift via distributionally robust optimization [Subbaswamy et al., 2021], use of the marginal sensitivity model for external shifts [Hatt et al., 2021], or variational characterizations of coherent risk measures [Jeong and Namkoong, 2020]. Methodologically, some of this work is similar to work in causal inference that studies unobserved confounding under the lens of robust optimization adversarial likelihood ratios over some ambiguity set [Kallus et al., 2018a, Kallus and Zhou, 2021, 2020, Dorn et al., 2021, Zhao et al., 2019, Bruns-Smith and Zhou, 2023, Yadlowsky et al., 2018, Tan, 2006]. This highlights the broad simultaneous interpretations of adversarial weight functions for handling unobserved confounding (in the generation of the data) in addition to robust adversarial covariate shifts (in the deployment of the predictor).
@@
Approaches based on multi-calibration inherently couple the specification of the (expanded) hypothesis class of multicalibrated predictors along with the specification of covariate shift functions, i.e. the boosting-type algorithm returns a predictor in the sum class of the original predictor and the classes of shifts. Approaches for robust covariate shift, to reduce the complexity of the adversary, require additional moment constraints satisfied by valid likelihood ratios, i.e specifying a sharp set $\mathcal{C}$ of only valid covariate shifts. In robust optimization-based approaches to covariate shift, the hypothesis class and class of weight functions can be independently varied. But for multi-calibration, restricting the auditor function classes also simultaneously reduces the functional complexity of the hypothesis class of predictors. Distributionally robust objectives are equivalent to variance regularization or control of the tail risk, which couples statistically more difficult control of tail behavior with the control of ambiguous shift functions. Another important point of difference is that the Bayes-optimal predictor satisfies the multi-accuracy criterion, while a Bayes-optimal predictor with heteroskedastic noise may not satisfy desiderata of uniform performance implied by distributional robustness. For example, [Duchi and Namkoong, 2018, Example 2] discusses the example of linear well-specified models where the distributionally robust predictor coincides with the Bayes-optimal predictor; but in cases of model misspecification/heteroskedastic noise, this may not be the case. We leave a finer-grained comparison for alternative work.
Finally, our discussion of the hybrid observational and randomized setting is more to highlight an “off-the-shelf” application of multi-accuracy, rather than the tightest analysis in this setting. Other works use more structure of this hybrid setting, or more heavily modify algorithms (i.e. learning shared representations) [Hatt et al., 2021, Yang et al., 2020, Kallus et al., 2018b]; analogous adaptations with multi-accuracy are interesting directions for future work. See also Bareinboim and Pearl [2013] for a survey on transportability and external validity, Colnet et al. [2020] for a survey on learning from observational and randomized data.

### v2

@@ 4.3 External validity: unknown deployment shift
However, the same properties of multi-accuracy also imply robustness to external shift. In this subsection, we indeed suppose Assumption 1, unconfoundedness. Recall that our goal was to control the predictive bias on a target covariate distribution $Q_{X}$, potentially unknown, $\lvert\operatorname{E}_{Q}[\left(\hat{\tau}(X)-\tau(X)\right)]\rvert$. Note that each of $\mu_{1},\mu_{0}$ are learned on a treatment conditional distribution, so we have that the valid likelihood ratio, which we denote $w_{t}(x)$, is defined as:
@@
Suppose Assumptions 1, 2 and 3. Let $\mathcal{C}$ denote a test function class for subgroup membership and $\mathcal{H}$ a test function class for likelihood ratios. Run Algorithm 1 for $\alpha$-multi-accurate postprocessing of the outcome models of a T-learner estimate. Then, for all target covariate distributions $Q$ such that the likelihood ratios $w_{1},w_{0}\in\mathcal{H},$
$\forall c(x)\in\mathcal{C},\;\lvert\operatorname{E}_{Q}[\{\tilde{\tau}(X)-(Y(1)-Y(0))\}c(X)]\rvert\leq 2\alpha.$
@@
Observe that although similar arguments apply, obtaining conditional guarantees for CATE estimation requires a richer test function class than for universal adaptability of the ATE alone. This illustrates that the case of learning CATE is indeed statistically harder than that of “universal adaptability” of the ATE. For CATE estimation, we need to choose a richer auditor function class than we would for ATE estimation.
4.4 Observational and randomized data (Setting 2)
(Meta)-Algorithm.
Identification and estimation of CATE.
Identification and estimation for the CATE follows by interpreting the auditing functions $c(X)\in\mathcal{C}$ as subpopulations. Achieving multi-accuracy on the RCT data hence identifies the CATE. That is, multi-accuracy assures us that
$\hat{\mu}_{t}(x)\leftarrow\arg\min_{g(x)\in\mathcal{G}}\operatorname{E}_{\text{obs}}[(g(X)-Y)^{2}\mid T=t],\text{ for }t\in\{0,1\}$
4:Return $\tilde{\tau}(x)=\tilde{\mu}_{1}(x)-\tilde{\mu}_{0}(x)$
Identification of target-independent CATE.
In complete analogy to the external shift setting, changing our interpretation of the target functions allows us to infer robustness to external shifts. Multi-accuracy ensures that, for all target covariate distributions $Q_{X}$ such that the likelihood ratios $w_{t}(x)\in\mathcal{H},\;t\in\{0,1\}$, running Algorithm 2 with auditor function class $\mathcal{F}=\mathcal{C}\times\mathcal{H}$ results in a multi-accurate and deployment-shift robust CATE estimate:
$\forall\;c(x)\in\mathcal{C},\;\lvert\operatorname{E}_{Q}[\{\tilde{\tau}(X)-(Y(1)-Y(0))\}c(X)]\rvert\leq 2\alpha.$
4.5 Extension to CATE pseudo-outcome regression
A natural question given our work on the $T$-learner is whether we can provide similar guarantees for an estimation-improved CATE learner, since the $T$-learner generally does not enjoy any improved estimation properties in causal inference. The causal inference and machine learning literature has developed many improved orthogonal/semiparametrically efficient procedures such as (but not limited to) the $R$-learner [Nie and Wager, 2020], $DR$-learner [Kennedy, 2023], or other machine-learning adaptations [Wager and Athey, 2018, Shalit et al., 2017].
Namely, some CATE estimation procedures give a pseudo-outcome $\psi(O;e,\mu)$, where $O$ denotes data tuples, i.e. $O=(X,T,Y)$, such that $\operatorname{E}[\psi(O;e,\mu)\mid X]=\tau(X)$. (It is designated as a pseudo-outcome because regressing upon it identifies the CATE or functional of interest, although it is not exactly an outcome itself). One such pseudo-outcome is the doubly-robust score. Pseudo-outcome regression of it as a CATE estimator was recently studied in Semenova and Chernozhukov [2021], Kennedy [2020].
@@ 4.5 Extension to CATE pseudo-outcome regression
Regressing upon pseudo-outcomes with favorable properties like orthogonality therefore confers these favorable properties to the estimated functional, such as improved statistical rates of convergence. Our arguments for external validity can naturally be extended for pseudo-outcome based CATE regression, so long as the pseudo-outcome’s conditional expectation is the CATE function.
We multi-accurately postprocess the pseudo-outcome regression step. That is, we learn $\tilde{\tau}$ such that:
@@
Multi-accurate DR-learner.
We give the algorithm for obtaining a multi-accurate $DR$-learner estimate in Algorithm 3. To summarize: we do need four folds of data $\left(\mathcal{D}_{1a},\mathcal{D}_{1b},\mathcal{D}_{2},\mathcal{D}_{3}\right)$; the first three for sample-splitting of the nuisance estimates and pseudo-outcome evaluation and the last for validation/calibration for MCBoost. Estimate the nuisance functions on the first two folds $\mathcal{D}_{1a},\mathcal{D}_{1b}$ and on $\mathcal{D}_{2},$ evaluate the pseudo-outcome value $\hat{\varphi}(O;\hat{e},\hat{\mu})$ and regress $\hat{\tau}(x)={\hat{\operatorname{E}}}_{n}[\hat{\varphi}(O;\hat{e},\hat{\mu})\mid X=x].$ Finally, we conduct post-processing via multi-accurate learning upon the $DR$-learner estimate $\hat{\tau}$, to obtain a multi-accurate $\tilde{\tau}$.
@@
5:Cross-fitting (optional)111 omitted for brevity because this may have to be further modified: original text reads, Repeat Step 1-2 twice, first using $\left(D_{1b}^{n},D_{2}^{n}\right)$ for nuisance training and $D_{1a}^{n}$ as the test sample, and then using $\left(D_{1a}^{n},D_{2}^{n}\right)$ for training and $D_{1b}^{n}$ as the test sample. Use the average of the resulting three estimators as a final estimate of $\tau$.
@@
Suppose Assumptions 2, 1 and 3. Suppose that $\tilde{\mu}_{t}(x)\leftarrow\textrm{MCBoost}(\hat{\mu}_{t},\alpha,\mathcal{F},\mathcal{D}_{post})$, over auditor function class $\mathcal{F}$ such that ¯$\overline{\mathcal{F}}_{t}\subseteq\mathcal{F}$, where ¯$\overline{\mathcal{F}}_{t}=\left\{1,\frac{\mathbb{I}[T=t]}{e_{t}(X)}\right\}\times\mathcal{C}\times\mathcal{H}$. That is, $\tilde{\mu}_{1}-\tilde{\mu}_{0}$ comprise an $\alpha$-multi-accurate $T$ learner. Then
$\Big{|}{\max_{ch\in\mathcal{C}\times\mathcal{H}}\operatorname{E}[\{{\varphi}(e,\tilde{\mu};X)-\tau(X)\}c(X)h(X)]-\max_{ch\in\mathcal{C}\times\mathcal{H}}\operatorname{E}[\{\tilde{\tau}(X)-\tau(X)\}c(X)h(X)]}\Big{|}\leq 2\alpha$
That is, the multi-accurate T-learner $\tilde{\mu}_{1}-\tilde{\mu}_{0}$ is, up to $2\alpha$ additive approximation error, a multi-accurate DR-learner with outcome model $\tilde{\mu}$, post-processed over the function class $\mathcal{C}\times\mathcal{H}.$
@@
$\max_{c\times h\in\{\mathcal{C}\times\mathcal{H}\}}\lvert\operatorname{E}[\{\{(Y-\tilde{\mu}_{1}(X))\frac{\mathbb{I}[T=1]}{e_{1}(X)}+(Y-\tilde{\mu}_{0}(X))\frac{\mathbb{I}[T=0]}{e_{0}(X)}\}+\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)\}-\tau(X)\}c(X)h(X)]\rvert$
$\qquad-\max_{c\times h\in\{\mathcal{C}\times\mathcal{H}\}}\lvert\operatorname{E}[\{\{\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)\}-\tau(X)\}c(X)h(X)]\rvert$
@@
Interestingly, concurrent with the preparation of this work, Bruns-Smith et al. [2023] study augmented balancing weights and find a certain target-independent property of the augmented estimator related to the universal adaptability of Kim et al. [2022]. Studying connections further would be an interesting direction for future work.
@@ 5.2 WHI data application
We next present a case study that draws on data from the Women’s Health Initiative (WHI) studies [Machens and Schmidt-Gollwitzer, 2003]. The WHI includes a large observational study and clinical trial data to investigate the effectiveness of hormone replacement therapy (HRT) in preventing the onset of chronic diseases. As the observational study has been suspected to suffer from various (unobserved) confounding phenomena [Kallus and Zhou, 2018], we study how utilizing both data sources in combination via multi-accuracy boosting compares to learning CATE from the observational or clinical trial data only. We focus on the effect of HRT treatment on systolic blood pressure and use age and ethnicity as covariates. Implementation details and results with an extended set of covariates is presented in the supplementary material D.3.
@@ 6 Related work: further discussion
A popular approach for handling unknown shifts is to enforce robustness against a family of covariate shifts (e.g. unknown shifts parametrized by unknown covariate shift functions) [Liu and Ziebart, 2014, Wen et al., 2014, Chen et al., 2016]. The goal is to find a robust hypothesis that maximizes the worst-case prediction risk (for example, squared error) evaluated with respect to unknown shifts within some class of covariate shifts. Parametrizations include distributionally robust optimization or linear basis functions. The work of Greenfeld and Shalit [2020] is motivated differently and penalizes with a Hilbert-Schmidt Independence Criterion loss; they show this implies some robustness to covariate shift. While most of this work is in the generic prediction setting, recent work also assesses ATE under covariate shift via distributionally robust optimization [Subbaswamy et al., 2021], use of the marginal sensitivity model for external shifts [Hatt et al., 2021], or variational characterizations of coherent risk measures [Jeong and Namkoong, 2020]. Methodologically, some of this work is similar to work in causal inference that studies unobserved confounding under the lens of robust optimization adversarial likelihood ratios over some ambiguity set [Kallus et al., 2018a, Kallus and Zhou, 2021, 2020, Dorn et al., 2021, Zhao et al., 2019, Bruns-Smith and Zhou, 2023, Yadlowsky et al., 2018, Tan, 2006, Bruns-Smith and Zhou, 2023]. This highlights the broad simultaneous interpretations of adversarial weight functions for handling unobserved confounding (in the generation of the data) in addition to robust adversarial covariate shifts (in the deployment of the predictor).
@@
Approaches based on multi-calibration or multi-accuracy inherently couple the specification of the (expanded) hypothesis class of multi-calibrated predictors along with the specification of covariate shift functions, i.e. the boosting-type algorithm returns a predictor in the sum class of the original predictor and the classes of shifts. Approaches for robust covariate shift, to reduce the complexity of the adversary, require additional moment constraints satisfied by valid likelihood ratios, i.e specifying a sharp set $\mathcal{C}$ of only valid covariate shifts. In robust optimization-based approaches to covariate shift, the hypothesis class and class of weight functions can be independently varied. But for multi-calibration, restricting the auditor function classes also simultaneously reduces the functional complexity of the hypothesis class of predictors. Distributionally robust objectives are equivalent to variance regularization or control of the tail risk, which couples statistically more difficult control of tail behavior with the control of ambiguous shift functions. Another important point of difference is that the Bayes-optimal predictor satisfies the multi-accuracy criterion, while a Bayes-optimal predictor with heteroskedastic noise may not satisfy desiderata of uniform performance implied by distributional robustness. For example, [Duchi and Namkoong, 2018, Example 2] discusses the example of linear well-specified models where the distributionally robust predictor coincides with the Bayes-optimal predictor; but in cases of model misspecification/heteroskedastic noise, this may not be the case. We leave a finer-grained comparison for alternative work.
Our discussion of the hybrid observational and randomized setting is more to highlight an “off-the-shelf” application of multi-accuracy, rather than the tightest analysis in this setting. Other works use more structure of this hybrid setting, or more heavily modify algorithms (i.e. learning shared representations) [Hatt et al., 2021, Yang et al., 2020, Kallus et al., 2018b]; analogous adaptations with multi-accuracy are interesting directions for future work. See also Bareinboim and Pearl [2013] for a survey on transportability and external validity, Colnet et al. [2020] for a survey on learning from observational and randomized data. In this literature, Wu and Yang [2023, 2022] among others consider reweighting towards shifted distributions. Other approaches include Cheng et al. [2023].